OpenStudy (anonymous):

Solve the equation: 6 10 ---- = ---- p-1 p+4 A. p = 1 B. p = 0.5 C. p = 7 D. p = 8.5

OpenStudy (texaschic101):

6/(p - 1) = 10/(p + 4) this is a proportion, so we can cross multiply 6(p + 4) = 10(p - 1) -- distribute 6p + 24 = 10p - 10 -- subtract 10p 6p - 10p + 24 = -10 -- subtract 24 6p - 10p = -10 - 24 -- combine like terms -4p = - 34 -- divide by -4 p = -34/-4 p = 8.5